value of cos(-660)*tan(510)
What is it value?
Please reply fast

Grade:11

2 Answers

Ajay
7 years ago
 
COS (-660) * Tan (510)
= Cos (660) * Tan (510) Since cos (-x) = Cos x
= Cos (360 + 300 ) * Tan (360 + 150 )
 = Cos 300 * Tan 150
= Cos (360 – 60) * Tan (180-30)
= Cos 60 * (-Tan 30 )
= ½ * (-1/sqrt 3)
= –1/2 * 1/ sqrt 3
raj
7 years ago
 
hello,
COS (-660) * Tan (510)
= Cos (660) * Tan (510) Since cos (-x) = Cos x
= Cos (360 + 300 ) * Tan (360 + 150 )
 = Cos 300 * Tan 150
= Cos (360 – 60) * Tan (180-30)
= Cos 60 * (-Tan 30 )
= ½ * (-1/sqrt 3)
= –1/2 * 1/ sqrt 3